OpenStudy (anonymous):

1. Joseph's auto sales last month were equal to two thirds of his sales this month plus 6. If he sold 9 cars this month, how many did he sell last month? A. 6 cars B. 21 cars C. 12 cars D. 9 cars

OpenStudy (travisbrown372):

2/3 of 9 = 6 6+6=12 so c.

OpenStudy (anonymous):

Ok but can u help me with 1 more please ??

OpenStudy (travisbrown372):

ok

OpenStudy (anonymous):

Solve the equation for x if ax + by = c x = (c - by)/a x = a(c - by) x = c - by - a x = ac - by

OpenStudy (travisbrown372):

x=(c-by)/a

OpenStudy (anonymous):

thank you

OpenStudy (travisbrown372):

yw